### v3

# Analysis of Failures and Risks in Deep Learning Model Converters: A Case Study in the ONNX Ecosystem

Abstract.
Software engineers develop, fine-tune, and deploy deep learning (DL) models using a variety of development frameworks and runtime environments. DL model converters move models between frameworks and to runtime environments. Conversion errors compromise model quality and disrupt deployment. However, the failure characteristics of DL model converters are unknown, adding risk when using DL interoperability technologies.
This paper analyzes failures in DL model converters. We survey software engineers about DL interoperability tools, use cases, and pain points (N=92). Then, we characterize failures in model converters associated with the main interoperability tool, ONNX (N=200 issues in PyTorch and TensorFlow). Finally, we formulate and test two hypotheses about structural causes for the failures we studied. We find that the node conversion stage of a model converter accounts for 75% of the defects and 33% of reported failure are related to semantically incorrect models. The cause of semantically incorrect models is elusive, but models with behaviour inconsistencies share operator sequences. Our results motivate future research on making DL interoperability software simpler to maintain, extend, and validate. Research into behavioural tolerances and architectural coverage metrics would be fruitful.
1. Introduction
Deep Learning (DL) achieves state-of-the-art performance in many domains (Grigorescu et al. 2020; Kim et al. 2019). Software engineers engage in many activities for deep learning, including developing, re-using, fine-tuning, and deploying DL models (Amershi et al. 2019; Han et al. 2021; Jiang et al. 2023c, a). They use tools at each stage: DL frameworks for development (e.g., PyTorch (Paszke et al. 2019)); DL model registries for re-use (e.g., HuggingFace (Face 2021)), and DL compilers for deployment platforms (e.g., TVM (Chen et al. 2018)). Preferably, these tools would be interoperable, so that DL models can move seamlessly from one to another. Model conversion errors disrupt engineering workflows or compromise the resulting models (Openja et al. 2022). High-quality model converters are crucial to the deep learning ecosystem.
Researchers have characterized failures in most of the DL ecosystem, but not in model converters. As depicted in Figure 1, previous works have considered DL development frameworks (Chen et al. 2022; Long and Chen 2022; Islam et al. 2019a; Pham et al. 2020; Tan et al. 2022) and DL deployment compilers and runtimes (Shen et al. 2021a; Guo et al. 2019). In contrast, prior research on DL model converters is limited to measuring conversions of 5 DL models (Openja et al. 2022). We lack systematic knowledge of failure symptoms, causes, and patterns in DL model converters.
In this work, we analyze failures of DL model converters. We first survey software engineers (N=108), focusing on their experiences with DL interoperability tools, their use cases, and failure modes encountered. Then we analyze failures in DL model converters to the ONNX IR (Open Neural Network eXchange’s Intermediate Representation), the most prominent interoperability target for DL models. We sample 200 closed GitHub issues (100 per converter) and determine failure symptoms, causes, and locations. Finally, we examine two possible root causes of model converter failures, testing hypotheses about specification updates and model types.
Our survey results show that ONNX is the most popular interoperability tool. It is primarily utilized for model deployment and framework conversion, with crashes and performance degradation being the most reported problems. Our failure analysis found that: common symptoms are crashes and incorrect model behaviors; common causes are incompatibility and type problems; and these issues tend to occur in a converter’s graph translation and graph optimization components. Most results were consistent between both systems examined as well as with prior studies. Finally, in our root-cause examination of why converters fail, we describe some model characteristics that are correlated with converter failures.
Our contributions are:
We survey 92 engineers and report common interoperability tools, use-cases, and pain point (section 4).
We analyze failures in two DL model converters: PyTorch and TensorFlow into ONNX. We taxonomize and measure the distribution of failure symptoms, causes, and locations (section 5).
We find that defective converter behavior is correlated with unusual models, but not with changes in the ONNX specification nor with the use of individual model layers or sequences (section 6).
Significance for software engineering: DL interoperability tools, especially model converters, underpin many deployment pipelines. We conducted the first systematic study of DL interoperability tools and model converters through a user survey and failure analysis. Understanding how and why these tools fail will help software engineers make informed judgments about their robustness.
2. Background and Related Work
Here we define DL model converters via the concept of interoperability, and discuss prior failure studies of DL ecosystem components.
2.1. DL Model Conversion as Interoperability
In the context of DL, interoperability focuses on model reuse and is defined as the ability of DL software to exchange DL models (i.e., deep neural networks/DNNs) (Liu et al. 2020). Wegner describes two patterns for interoperability: creating pairwise mappings between systems; and introducing a common intermediary understood by all participants (Wegner 1996). These patterns are illustrated in Figure 1 in the context of DL. The two patterns trade-off customizability for scalability.
Two kinds of DL systems interoperate on DL models: frameworks for DL model development, and runtimes for DL model deployment. In the downward path of Figure 1, pairwise mappings occur in DL compilers such as TVM (Chen et al. 2018), which map from framework representations into internal compiler representations for each supported framework. Along the rightward path, common intermediaries such as ONNX (ONN 2019) and MMDnn (Liu et al. 2020) give standard representations for DL models. In this pattern, each framework and runtime has a one- or two-way adapter to the common intermediary.
DL Model Converters: Model converters fill a purpose similar to compiler front-ends (Li et al. 2020). They transform a model from a DL framework into a high-level IR representing the model’s computations and control flow. Graph-level optimizations are applied before further conversion to a low-level IR for hardware optimization and code generation. Table 1 summarizes a typical design.
| Component | Definition |
| Load Model | Framework representation ONNX graph. Tracing used for dynamic graphs (e.g., PyTorch). |
| Node conversion | Graph nodes replaced by ONNX equivalents. |
| Optimization | Nodes (e.g., operator fusion), dataflows (e.g., DCE). |
| Export | Model serialized into protocol buffer (protobuf). |
| Validate | Syntactic checks (compliance with spec) and semantic checks (behavioral changes). |
Figure 2 illustrates a model converted from PyTorch to ONNX. Conversion is challenging, as noted by AirBus (ONNX 2023; Gauffriau and Pagetti 2023) and others (Consortium 2023), because it maps between graphs expressed with different operators and different semantics. Model conversion can produce models that are incompatible with runtimes or have different behaviours. For a compatibility issue, in PyTorch #78721 a converted ONNX model had a type mismatch (vbogach 2022). For a behavioral issue, in PyTorch #74732 a converted ONNX model’s prediction changed (YsYusaito 2022).
After conversion to an IR, models can be rendered back to DL frameworks or deployed to hardware. Framework-to-framework conversion can bypass issues in model reengineering (Jiang et al. 2023b). Deployment from an IR allows DL runtimes to optimize against the IR rather than the many DL frameworks.
2.2. Failure Studies of DL Components
Software engineering failures inform development and maintenance (Petroski et al. 1994; Anandayuvaraj and Davis 2022; Amusuo et al. 2022a). Previous failure analyses of DL interoperability software focused on the “Development” and “Runtime” components of Figure 1. Chen et al. studied 800 defects from 4 DL frameworks to obtain testing guidelines (Chen et al. 2022). Shen et al. studied DL compiler defects (Shen et al. 2021a), and others have tested DL compilers (Liu et al. 2023; Xiao et al. 2022). Some of this work has incidentally examined interoperability failures (Liu et al. 2023; Morovati et al. 2024); for example, Shen et al.’s study of DL compilers included “front ends” offering interoperability through pairwise (DL framework specific) and common intermediary (ONNX) model loaders (Shen et al. 2021a);
The two prior failure studies of DL model converters focused on testing and fault localization. Openja et al. converted 5 popular DL models (Openja et al. 2022). They considered only the current state of model converters, not past failures. Louloudakis et al. studied behavioral issues resulting from framework-to-framework conversion (Louloudakis et al. 2023a). They found failures in 10 out of 36 conversions. They created a fault localization and repair pipeline to localize and fix discrepancies (Louloudakis et al. 2023b).
In light of this literature, the main contribution of our work is the first systematic analysis of failures in DL model converters. As a conceptual contribution, we frame DL model converters as a class of interoperability software. We consider how and why they fail.
3. Research Questions & Study Design
Figure 3 shows our RQs (section 3.1) and study design (section 3.2), detailed next.
3.1. Research Questions
We summarize the related work in section 2 as follows: The DL ecosystem is growing more complex (section 2.1), motivating a shift to common intermediaries such as ONNX. Analyzing failures in this emerging pattern will inform ecosystem participants of risks and opportunities for improvement. Prior work examined failures in DL frameworks and runtimes, but has paid little attention to interoperability (section 2.2).
Our study fills this knowledge gap by analyzing failures in DL interoperability software. We specifically focus on ONNX, the leading DL interoperability framework. Our research proceeds in three themes: (1) why and how engineers use interoperability tools; (2) a failure analysis of the most popular interoperability tool; and (3) evaluating hypotheses about the root causes of failures.
Theme 1: Interoperability User Survey (section 4)
RQ1
How and why do engineers use interoperability tools?
Theme 2: Failure Analysis (section 5)
RQ2
What are the failure characteristics in DL interoperability software — symptoms, causes, and locations?
Theme 3: Hypotheses on ONNX Failure Causes (section 6)
RQ3
Does ONNX evolution affect converter failure rates?
RQ4
Do model types affect converter failure rates?
3.2. Study Design
Figure 3 relates research questions to methods. For RQ1 we use a user survey (section 4). The survey results inform the remainder of the work. For RQ2 (section 5) and RQ3-5 (section 6) we apply methods from mining software repositories and software testing. We analyze GitHub issues (RQ2), correlate issue frequency to ONNX versions (RQ3), and evaluate ONNX on a range of model types (RQ4).
4. Theme 1: Interoperability Study
To understand the use of DL interoperability, we surveyed DL practitioners on DL development and deployment practices. We selected the survey methodology to gather insights across diverse practices and experiences (Easterbrook et al. 2008). Method is in section 4.1, results in section 4.2.
4.1. Methods
4.1.1. Instrument Design
We followed Kitchenham & Pfleeger’s guidelines to develop our survey instrument (Kitchenham and Pfleeger 2008). Table 2 illustrates the result. We asked general questions about DL interoperability, and specific questions about ONNX (interview data (Shankar et al. 2024) and GitHub suggest it is most popular). For higher response rate, the survey was short and mostly closed-ended. We iterated internally, piloting with 3 external participants to check instrument clarity.
We sent the survey out in batches to allow for iteration. We examined the results halfway to determine whether the instrument aligned with a larger sample of users’ behavior. We observed an unexpectedly high incidence of one use case, and so in the second half we added an open-ended question to clarify this use case.
| Topic | Example questions |
| Demographics | (1) How long have you worked on ML/DL projects? (2) What deployment environments are targeted? |
| Interoperability tool usage | (1) What do you consider when choosing between deploying from a DL framework vs. via a tool like ONNX? (2) Do you use ONNX as part of your model development and deployment process? |
| Use cases | (1) For what purpose do you use interoperability tools? (2*) If you use ONNX for framework-to-framework conversion, please describe your use case further? |
| Failure modes (ONNX-specific) | (1) Do you commonly encounter problems while working with ONNX models? (2) If you encounter such problems, how do you address them? |
4.1.2. Population and Sampling
With approval from our Institutional Review Board (IRB), survey respondents were recruited from Hugging Face users. The Hugging Face ecosystem is the primary location for deep neural network-based model development and re-use (Jiang et al. 2023d), so its users are a suitable population for questions about DL interoperability tools. To increase the likelihood of responses from experienced software engineers, we collected email addresses from users with PRO accounts (i.e., paid), and from accounts in organizations marked as company, community, and non-profit (excluding types such as education). Participants received $10 gift cards.
We targeted a confidence level of 90% with a 10% margin of error. With an estimated Hugging Face user population of 1.2 million (Forbes 2024; Gargi 2024), a sample size of 69 respondents was needed. We sent out surveys in batches until reaching our desired sample size. In total, we distributed our survey to 228 PRO users and 1,985 organization members. We received a total of 92 valid responses (4% response rate). All questions were optional to improve the response rate, so we do not have responses from all subjects on all questions.
Table 3 shows respondent demographic information.
| Type | Break-down |
| ML Exp. | 1 yr. (8) ; 1-2 yr. (17) ; 3-5 yr. (32) ; 5 yr. (32) |
| SE Exp. | 1 yr. (4) ; 1-2 yr. (12) ; 3-5 yr. (20) ; 5 yr. (53) |
| Org. Size | Small, employees (48) ; Medium, employees (17) ; Large, employees (24) |
| Deployment Env. | Web application (59) ; Cloud and data center (52) ; Desktop application (19) ; Mobile (14) ; IoT/embedded systems (14) ; Other (4) |
4.1.3. Analysis
Most questions were closed-ended (multiple-choice, checkbox). Qualitative analysis was needed for 3 open-ended questions, about use cases, model deployments, and interoperability problems. Two authors reviewed the data and agreed that the responses to these questions were short and did not involve much subjectivity. Therefore, all data were analyzed by one author.
4.2. Results
4.2.1. Interoperability Tool Usage
The majority of respondents utilized the PyTorch or TensorFlow frameworks for model development. Most of the respondents utilized PyTorch (89%). TensorFlow is the second most used framework (37%). JAX, MLX, and other frameworks made up 20% responses. The data show that the respondents often use multiple frameworks or have moved from one framework to another (e.g., moving from TensorFlow to PyTorch), perhaps motivating their use of interoperability tools.
Approximately 42% (39/92) of respondents reported using ONNX. Only 15 respondents used other interoperability tools such as MMdnn and NNEF. The rest (41%) do not use interoperability tools. In our survey, although we did not gather extensive data on participants’ specific roles, insights from their deployment considerations suggest they are primarily model developers. They either lack model deployment responsibilities or deploy the models directly.
4.2.2. Use Cases
Interoperability tools are used equally to interoperate between frameworks and to deploy models. 69% (37/54) of respondents report using model conversion for deployment of DL models. Whereas 52% (28/54) of respondents report using framework-to-framework conversion (e.g., PyTorch to TensorFlow).
For the respondents who use interoperability tools for framework-to-framework conversion, most of them use the tools to “integrate the models in non-Python code”. For example, one respondent wants to “interact with the base CUDA system”. Another respondent “exports [pre-trained models] to ONNX for import into Axon, an Elixir/Nx based DL framework that often [lacks] native pre-trained models”.
Table 4 presents the deployment considerations of practitioners when using interoperability tools. Many practitioners report deploying directly from frameworks when they want easy deployment or expect to update models often. The broader organization’s practices may play a role: “if I have control over the entire training to deployment pipeline, it is easier to use PyTorch exports”.
| Theme | # Participants (%) |
| Simplicity | 8 / 32 (25%) |
| Deployment requirements | 7 / 32 (22%) |
| Inference speed | 5 / 32 (16%) |
| Portability | 5 / 32 (16%) |
| Other (e.g., maintenance, stability) | 2 / 32 (6%) |
| Deploy directly from the DL frameworks | 13 / 32 (35%) |
Interoperability tools are preferred when respondents care about performance gains, portability, language compatibility, or if they require exotic configurations. Examples were: “[PyTorch in browser is] too big”, and “[ONNX offers] portability over performance”.
4.2.3. Failure Modes
We ask practitioners about their experiences using ONNX. 59% (32/54) of practitioners report encountering problems when using ONNX, while 41% (22/54) report that they do not commonly encounter problems with ONNX models. Crashes (19/32, 59%) and performance differences (19/32, 59%) are the most prominent. Often models do not convert, or there are performance differences between the original and converted models. There are also some other problems (6/32, 19%) mentioned by the respondents. Examples included “ONNX doesn’t support Fourier layers” and “driver problems during…deployment on user machines”.
Table 5 presents practitioners’ strategies for resolving issues with the ONNX converter. Many (44%) turn to community resources, including GitHub issues and Stack Overflow posts, for help. One-third (31%) verify the conversion process through testing, while 19% explore solutions by changing ONNX versions. Some had experienced no good solution, resolving ONNX issues “case-by-case”.
| Theme | # Participants (%) |
| Seeking help from the community | 7 / 16 (44%) |
| Test with executions | 5 / 16 (31%) |
| Version changes | 3 / 16 (19%) |
| Other (i.e., documentation, config) | 3 / 16 (19%) |
5. Theme 2: Failure Analysis (of ONNX)
The second theme of this work is a failure analysis of deep learning interoperability software, specifically, deep learning model converters. We apply the method of Failure Analysis (Petroski et al. 1994; Leveson 1995; Amusuo et al. 2022b; Anandayuvaraj and Davis 2022), which characterizes and reports the distributions of past failures to revise engineering methodologies and prioritize research targets.
Many model converters have been proposed, but our survey data show which converters are of practical interest (section 4.2.1): those for the ONNX framework (half of respondents use it); and specifically those associated with the PyTorch and TensorFlow model development frameworks (the two most common).
About ONNX: The ONNX (Open Neural Network eXchange) specification provides a common representation that DL frameworks and runtimes use to represent DL models (ONNX 2022). The ONNX specification has three main components: (1) A definition of a computational graph; (2) definitions of standard data types; and (3) definitions of built-in operators such as ArgMax (ONNX 2022). The ONNX specification changes regularly to keep up with DL framework evolution (Shridhar et al. 2020). These changes consist of adding new types and operators, and updating the behavior of existing operators. These changes are versioned within operator sets. As of 2023 there have been 18 operator sets, totaling 149 additions and 217 updates to the available operators.
5.1. Methods
5.1.1. Data Selection
Figure 4 depicts our data selection method. We describe the five main stages and rationales next.
(1) Repositories: For the reasons noted above, we studied the DL model converters from PyTorch and TensorFlow into the ONNX IR (torch.onnx and tf2onnx, respectively). We note that among ONNX model converters, those for PyTorch and TensorFlow have the most failure data available on GitHub (Table 6).
| Project | Input | Stars | Forks | Closed Issues |
| torch.onnx | PyTorch, Caffe2 | 77,429* | 20,938 | 1,286 (959)* |
| tf2onnx | TensorFlow, Keras | 2,200 | 425 | 848 (792) |
| Paddle2ONNX | PaddlePaddle | 637 | 145 | 192 |
| sklearn-onnx | Sci-kit Learn | 506 | 94 | 328 |
| onnx2torch | ONNX | 370 | 35 | 32 |
| onnx-tf | ONNX | 291 | 25 | 129 |
| onnx-coreml | ONNX | 385 | 79 | archived |
Using the PyTorch and TensorFlow converters covers two relevant differences within DL and interoperability. Within DL, these converters include both common representations of computational graphs: static (TensorFlow) and dynamic (PyTorch) (Hattori et al. 2020; Baydin et al. 2018). Within interoperability, these converters include both kinds of converter owners. Converters are naturally owned either by the upstream producer of the data or the downstream consumer of the data. In our case, torch.onnx gives an example of upstream ownership (it is owned by the PyTorch engineers), while tf2onnx gives an example of downstream ownership (it is owned by the ONNX engineers).
(2) Filters for relevance: For each repository, we collected closed GitHub issues related to ONNX conversion. For torch.onnx, the PyTorch repository marks ONNX converter issues with the label module: onnx. We collect all closed issues with this label, yielding the search filter: is:issue label:“module: onnx” is:closed. For tf2onnx, all issues are relevant. We remove issues labeled “pending on user response” and “question” because these issues may not be failures. This yields in 959 issues in torch.onnx (from 20,782 issues, not all ONNX-related) and 649 issues in tf2onnx (from 792 issues).
(3) Filters for data availability: Following prior work (Islam et al. 2019b; Franco et al. 2017; Garcia et al. 2020), we subsequently filtered for GitHub issues that contained enough information for failure analysis. We filter issues for sufficient information (e.g., the issue is resolved with a commit and pull request). This filtering is conducted upon the timeline events for each GitHub issue and the filtering criteria are given in Figure 4. This filter yields 327 issues in torch.onnx and 242 issues in tf2onnx.
(4) Filter Validation: We piloted filters to ensure they captured relevant issues (recall) but not irrelevant issues (precision). We hand-labelled issues per repository prior to filtering, applied the filter, and measured recall and precision. If recall and precision we consider our filter to acceptably balance manual work against bias. For torch.onnx, we measured recall of and precision of . For tf2onnx, we measured recall of and precision of .
5.1.2. Data Analysis
We describe converter failures in 3 dimensions: location, symptoms, and causes.
Location: We map failures to the converter stages in Table 1: Load Model, Node Convert, Optimization, Export, and Validate.
Symptoms & Causes: We adapted taxonomies from Shen et al. (Shen et al. 2021a), the closest related work. As that work concerns DL compilers rather than model converters, we trialed the taxonomies on a sample of 30 randomly chosen issues from each repository. Two researchers classified symptoms and causes. Computing inter-rater reliability via the Kappa coefficient (Cohen 1960), we found =0.90 (causes) and =0.95 (symptoms), i.e., “strong agreement” (McHugh 2012) for both taxonomies.
The taxonomy we used for symptoms is: crash, wrong model, bad performance, build failure, and unknown/other. In a crash, conversion errors out. A wrong model is syntactically sound but semantically incorrect. Bad performance indicates unexpectedly high time or memory cost, e.g., worse than on a previous version. Build failure means the user could not install the converter.
The taxonomy of failure causes is more complex — see Table 7.
| Cause | Definition |
| Incompatibility - Internal | API compatib. issues in model converters. |
| Incompatibility - External | Compatibility issues with third-party libraries (e.g., ONNX, TensorFlow). |
| Type Problem - Node | Issues related to atomic DL operators in model conversion. |
| Type Problem - Tensor | Issues related to the types of tensors. |
| Type Problem - Conventional | Issues with types of conventional variables in software systems. |
| Tensor Shape Problem | Input/output tensor shape issues. |
| Alg. Error - Optimization Error | Issues in model converter optimizations. |
| Alg. Error - Tracing Issue | Issues tracing dynamic models in computational graphs. |
| Alg. Error - Non-optim. Code Logic | Bugs outside DL compiler optimizations. |
| Testing | Test errors, incl. flaky or missing tests. |
Interrater Agreement: In our pilot study of applying Shen et al.’s taxonomies, we observed strong agreement between raters. Therefore, we relied on a single rating (one rater analyzing all issues) of the 200 sampled issues. To assess the risk of taxonomic drift (Wicks 2017), a second rater analyzed one tranche of randomly selected failures. In those 20 samples, the two raters had perfect agreement/=1.0. We conclude the single-rater analysis was sound.
5.2. RQ2: The Characteristics of Failures
5.2.1. Failure Locations
tf2onnx and torch.onnx have similar failure location distributions (Table 8). The most common location of failures is the Node Conversion stage: 148/200 failures occur in this stage, with a similar proportion in tf2onnx (70%) and torch.onnx (78%). The Graph Optimization stage is the second most common location of failures, with 19 in total. The distribution of failures in this stage differs between converters, with 3x more in tf2onnx.
| Location | TF | PT | Total |
| Load Model | 5 | 6 | 11 (6%) |
| Node Conversion | 70 | 78 | 148 (74%) |
| (Graph) Optimization | 14 | 5 | 19 (10%) |
| Export (Protobuf) | 1 | 0 | 1 (1%) |
| Validation | 0 | 3 | 3 (2%) |
| (Not Distinguishable) | 10 | 8 | 18 (9%) |
| Total | 100 | 100 | 200 (100%) |
5.2.2. Failure Symptoms
The distributions of symptoms for both tf2onnx and PyTorch are similar — see Table 9. The most common failure symptoms are Crash and Wrong Model, comprising 85% of failures in each converter. The Bad Performance and Build Failure make up around 5% of failure symptoms.
Our results show the similarity of the failure symptom distribution between DL model converters (our work) and DL compilers (Shen et al. (Shen et al. 2021a)). In the last two columns of Table 9, we see the Crash and Wrong Model symptoms make up the majority of symptoms across both DL converters and DL compiler front-ends. The Wrong Model and Crash symptoms appear characteristic of interoperability in this context, regardless of implementation.
| Symptom | TF | PT | Total | DL Comp. (Shen et al. 2021a) |
| Crash | 50 | 62 | 112 (56%) | 226 (63%) |
| Wrong Model | 35 | 30 | 65 (33%) | 100 (28%) |
| Build Failure | 3 | 2 | 5 (3%) | 3 (1%) |
| Bad Performance | 2 | 1 | 3 (2%) | 6 (2%) |
| Hang | 0 | 0 | 0 (0%) | 4 (1%) |
| Unreported | 10 | 5 | 15 (8%) | 20 (6%) |
| Total | 100 | 100 | 200 (100%) | 359 (100%) |
5.2.3. Failure Causes
We report the joint distribution of failure causes by symptom in Table 10. The most common failure causes are the Incompatibility and Type Problem, with more than 50% of failures exhibiting these causes for both converters. Algorithmic Errors and Shape Problems each contribute 20% of cases. By symptom, crashes were caused by incompatibility and type problems, while wrong models were caused by type problems and algorithmic errors.
For most causes, we see a similar distribution in each studied converter. For Algorithmic Error, we observe that tf2onnx has three times as many as torch.onnx. The disparity in Algorithmic Errors varies by the subclass, with the majority in torch.onnx being related to the loading of models namely tracing. An example is PyTorch #84092: trace not working on autocast. In contrast, in tf2onnx the majority are related to optimizations, such as tf2onnx #226 (incorrect reshape) and tf2onnx #1719 (incorrect folding). We conjecture that this difference may have a deeper cause: recall that the PyTorch converter is owned by the PyTorch engineers, while the TensorFlow converter is owned by the ONNX team (section 5). This difference in ownership may reduce the team’s understanding of the implications of optimizations, leading to worse outcomes in TensorFlow.
| Cause Symptom | Crash | Wrong Model | Other | Total | ||||
| TF | PT | TF | PT | TF | PT | TF | PT | |
| Incompatibility | 19 | 28 | 4 | 3 | 2 | 1 | 25 | 32 |
| Type Problem | 8 | 14 | 17 | 13 | 0 | 2 | 25 | 29 |
| Algorithmic Error | 4 | 3 | 10 | 3 | 4 | 0 | 18 | 6 |
| Shape Problem | 5 | 4 | 4 | 7 | 0 | 1 | 9 | 12 |
| API Misuse | 6 | 5 | 0 | 1 | 0 | 0 | 6 | 6 |
| Other | 8 | 8 | 0 | 3 | 9 | 4 | 17 | 15 |
| Total | 50 | 62 | 35 | 30 | 15 | 9 | 100 | 100 |
6. Theme 3: Investigating Deeper Causes
In the final theme of this work, we investigate the possibility that ONNX converter errors have a shared structural cause. By this, we mean a latent cause beyond the code-level causes in the taxonomy of Table 7. If so, ONNX users could use this factor to better assess their risk for the associated failure modes. Based on our survey data (section 4) and failure analysis (section 5), we formulate and test two hypotheses of structural failure causes. The RQs and hypotheses are:
Does ONNX evolution affect converter failure rates? RQ3: Does ONNX evolution affect converter failure rates? We hypothesize H: Changes in ONNX operator sets are correlated with increased defects. This hypothesis is based in data from our survey and failure analysis. In the survey, 19% of respondents reported changing ONNX versions as a possible solution (section 4, Table 5). This suggests that conversion defects (not just compatibility issues) may be localized by version. In the failure analysis, we found that crashes were largely due to Incompatibilities or Type Problems (section 5.2, Table 10). These failure modes relate to API compatibility and conversion correctness, which can be affected by changes in the ONNX opset.
RQ4: Do model types affect converter failure rates? We hypothesize H: Failures are caused by model structure, i.e., models with particular layers (node) or layer sequences are more prone to defects. This hypothesis is primarily based on our failure analysis. The majority of failures occurred during Node Conversion (Table 8), indicating that nodes and node sequences may be problematic. Survey data also had a relevant anecdote: one respondent solved conversion issues by re-implementing models using pure tensor operations, implying that conversion may fail due to the presence of certain operations.
In the remainder of this section, we describe the method and result for testing each hypothesis. Ultimately, we find evidence to support H. This result will help guide future validation efforts (section 7.1).
6.1. RQ3: Does ONNX evolution lead to failures?
Here we describe the method and results for testing H: That changes in ONNX operator sets are correlated with increased defects. The expected correlated failures in the model converters are incompatibility and type problems (defined in Table 7).
6.1.1. Method
We test the hypothesis by checking whether larger changes in the ONNX specification correlate with greater incidence of failures in DL model converters. To obtain the size of changes to the ONNX specification, we count the number of operator changes per release in the ONNX release notes (Figure 5). We approximate the number of failures per release using the GitHub issue creation times of the failures sampled for RQ2. We then attribute each issue to the nearest previous ONNX release.
We note four ways in which this measurement approximates: (1) DL model converters lag behind ONNX releases (this might cause a failure to be mis-attributed to another reuse, i.e., offset in time) ; (2) Failures might be in any ONNX available release, not just the most recent (possibly inflating the failure rate of a given release); (3) our failure analysis data were randomly sampled, possibly under-sampling certain time windows (though we note that our sample comprises 30-50% of qualifying issues); (4) as noted in the user survey (section 4.2.3), users may simply revert to a previous operator set without opening a bug report. For items 1 and 2, issues do not reliably include ONNX versions, so no more accurate data is available. For item 3, we use the data sampled during RQ2 so that we can check incidences by cause (incompatibility and type problems).
We test the hypothesis qualitatively and quantitatively. Qualitatively, we inspect a visualization of the hypothesized trend. We also measure the relationship, assessing the correlation in the number of changes in an ONNX release and the number of failures between its release and the next. We use the Spearman correlation, which is a commonly-used and robust metric for measuring a monotonic relationship between two variables (Fenton and Bieman 2014).
6.1.2. Results
Figure 6 depicts the hypothesized correlation between ONNX releases and model converter failures from section 5. After larger ONNX releases (Figure 5) we expect more failures (Figure 6).
Qualitatively, in Figure 6 we see no discernible increase in the number of failures following larger ONNX updates. Quantitatively (Spearman), results are similar. The test yields a weak positive correlation (). Similarly, Incompatibility and Type Problems are weakly positively correlated (). In the test, we discarded the first release (converters may be unstable) and the most recent release (insufficient data). We also merged releases 2–7 because they are too close together (released within a 1-month period).
Given the weak correlations, we do not find evidence to support H. There is a fairly steady rate of defects in ONNX model converters, whether the associated ONNX release is large or small.
6.2. RQ4: Do model types affect failure rates?
For this RQ we analyze models at two scales, macro and micro. In our macro scale analysis (section 6.2.1, section 6.2.2), we study entire models (all layers). For our micro scale analysis (section 6.2.3, section 6.2.4) we study individual layer and sequence patterns from the macro analysis.
6.2.1. Method for macro analysis
At the macro scale, we test DL converters using two types of models and then analyze the failure-inducing inputs. We used a differential testing approach (Figure 7). Differential tests need (1) inputs, and (2) a difference criterion (McKeeman 1998).
(1) Inputs: For inputs, we converted both real models and synthetic models. Real models contain input patterns that the converter expects. Synthetic models are more diverse and should exercise edge cases on the converter. Real and synthetic models came from:
Real Models: We used the HFTorrent dataset from Jiang et al. (Jiang et al. 2023d). At time of experiment, this was the largest available set of real-world DL models, containing 63,182 pre-trained models collected from the HuggingFace model registry. We filtered for the 1,761 models with both PyTorch and TensorFlow versions, allowing comparison between the two converters on similar inputs. These models represent 112 distinct full architectures (58 backbones).
Synthetic Models: We systematically generated synthetic DNNs for conversion, using the NNSmith tool (Liu et al. 2023). NNSmith generates random valid TensorFlow and PyTorch DNNs that use operators supported by ONNX. We added a parameter to NNSmith to generate DNNs of a given size (# nodes). We then systematically generated DNNs containing between 15-100 nodes in increments of 5. For each node count, we generated DNNs until two conditions were met: (1) All NNSmith-supported operators appeared at least once in the family; and (2) 100 distinct models were created.
In initial tests, we found the resulting models were often convertible yet unsupported by ONNX RunTime. To address common errors, we constrained the NNSmith synthesizer. Specifically, we removed 1 operator out of 71 for torch.onnx and 3 out of 56 for tf2onnx. Additionally, we confined the tensor data type to float32.
There are other DL model generation approaches, e.g., MUFFIN (Gu et al. 2022), COMET (Li et al. 2023), and LEMON (Wang et al. 2020). These tools generate inputs by mutating seed models, typically real models. We viewed HFTorrent as a sufficient source of real models, and synthesized with NNSmith as a complementary approach.
(2) Difference criterion: For each model, we attempt to convert it to ONNX using its respective converter. If conversion succeeds, we load it into the ONNX RunTime. For models that load successfully, we perform inference on both the original model (using PyTorch or TensorFlow) and the converted model (using ONNX Runtime). For real models, we use the inputs provided by the model owners to test the model. (These are also known as “dummy inputs” on the HuggingFace platform.) For synthetic models, we use 100 randomly generated inputs matching the model’s input shape.
In both cases, to measure model misbehavior we used a simple community-accepted approach: the distance measure and threshold used by the PyTorch exporter. This rule is that between the original and ONNX-converted models, in the inference result tensors, the maximum absolute element-wise difference should be (PyTorch 2023).
We considered alternatives to measure behavioral differences. A more general notion of tolerance or “acceptable error” could be used, but the choice of tolerance is not well established — the PyTorch verification tool uses , NNSmith uses (Liu et al. 2023). Openja et al. (Openja et al. 2022) proposed using model accuracy and robustness, but this method requires training each model on a suitable dataset. Training substantially increases the cost of the measurement, and the dataset requirement limits the types of models that can be generated.
6.2.2. Results for macro analysis
Real Models: Real models exhibited few converter failures and little incorrect model behavior. Most issues occurred within the HuggingFace-specific converter. Converter failures occurred for only 85/3,522 models (2 %). Once models reached torch.onnx and tf2onnx, over 90 % could convert and had equivalent behavior in ONNX Runtime.
One interesting form of failure we observed was models with identical architectures but varying conversion issues. For example, in tf2onnx, various checkpoints of the t5 model had 65 HuggingFace conversion errors, 56 unsuccessful tf2onnx conversions, 1 unsuccessful ONNX Runtime load, 0 incorrect outputs, and 183 successes. We could not determine the cause(s) of this instability.
Synthetic Models: Synthetic models exhibit more difficulties and often show incorrect model behaviour.
Crashes: Synthetic models exhibit difficulties when converting to ONNX and loading converted models to the ONNX Runtime. We find 802/3,544 (23%) Unsuccessful Conversions, almost entirely in synthetic models on tf2onnx. These crashes correspond to 4 unique crash locations. Of the 2,742/3,544 (77%) synthetic models that successfully convert, only 1,244/2,742 (45%) successfully load into ONNX Runtime. We observed 11 unique ONNX Runtime errors, of which 6 appear to correspond to open GitHub issues (vbogach 2022; josephrocca 2021; BowenBao 2023; rafaelagrc 2022; fatcat z 2022; 12sf12 2022). We disclosed the rest to the engineering teams. For most disclosures, we have not yet received a response. For one disclosure, which described a model that causes the libc++ abi to terminate unexpectedly during inference with ONNX Runtime, the response was that it had been (unintentionally) fixed in the development branch. It still affects older releases.
Behavioural differences: We observed a large fraction of behavioural differences (incorrect output) with synthetic models. Compared to real models, which had 20 instances, synthetic models had 320 instances where the inference results exceeded the threshold. The majority of these instances were observed in the tf2onnx converter. For both converters, we disclosed such instances to the respective engineering teams (we have not heard a response yet). A summary of disclosed issues can be found in the artifact.
| Outcome | tf2onnx | torch.onnx | ||
| Real | Syn. | Real | Syn. | |
| Start: Total models | 1,761 | 1,820 | 1,761 | 1,724 |
| Conv. Fail (HF) | 456 (26%) | N/A | 342 (20%) | N/A |
| Conv. Fail | 65 (4%) | 800 (44%) | 20 (2%) | 2 (1%) |
| ORT load Fail | 19 (1%) | 757 (42%) | 27 (2%) | 741 (42%) |
| Mismatch | 9 (1%) | 220 (12%) | 11 (1%) | 100 (6%) |
| Successful | 1,212 (68%) | 43 (2%) | 1,361 (75%) | 881 (51%) |
6.2.3. Method for micro analysis
To investigate the causes of the conversion failures we observed, we analyze mismatching models in terms of the operators used. We examine the individual operators and the sequences of operators. We compare these to non-failing models (for trends) as well as to the models in the converter test suites (for gaps in testing). Test suite models are collected from converters’ CI/CD pipelines. Figure 8 illustrates our approach.
For operator types, we measure the operators present in each converted model, out of the set of available operators in the ONNX opset. For operator sequences, we extract the simple paths through each model. To illustrate, the ONNX model in Figure 2 shows three operators (ReduceMax_0, ArgMax, and ReduceMax) and two simple linear paths (ReduceMax_0ArgMax and ReduceMax_0 ReduceMax). Models with similar architectures may have simple paths that are largely identical, inflating the number of shared sequences. For example, the sequences aaaabd and aaaabc are identical except for the last element. To deal with cases we further reduce the common sequences to the smallest shared subsequences, i.e., we recursively find the longest common subsequences for our sequences until we cannot find any smaller subsequences (Cormen et al. 2022). After reduction we find that sequences are between 3 to 5 operators long.
Evaluation Criteria for H: To evaluate the impact of operator types, we compare the operators of mismatched and correct models. If failing and correct models use different operators, we conclude that specific operators may cause failures. To evaluate the impact of operator sequences, we identify shared operator sequences belonging only to mismatched models, and their frequency. If mismatching models often share common simple paths of operators, that will support H. H would be further strengthened if operator sequences are rarely shared with correct models.
6.2.4. Results for micro analysis
We focused on synthetic models that converted successfully but had mismatching behavior. There were 330 such models across TensorFlow and PyTorch. We compared them to correct models and models from converter test suites.
Incorrect conversions cannot be explained solely by operator types. Of the 154 ONNX operators, mismatched models contain 58 (torch.onnx) and 54 (tf2onnx) unique operators while correct models contain 59 (torch.onnx) and 52 (tf2onnx) unique operators. All but 1-2 operators are shared, indicating that for synthetic models, operator types do not predict mismatch.
In contrast, our analysis of shared operator sequences supports hypothesis H. As shown in Table 12 an overwhelming majority of mismatched models share unusual operator sequences. of the mismatching models tested share at least one operator sequence that never occurred in the correct models. Further, we assess how often the sequences are shared by many models. We find that for torch.onnx, only 1 of the 131 sequences is shared by more than 3 models, and for tf2onnx no sequence is shared by more than 11 models. This indicates the failing models will often contain common operator patterns, suggesting families of sequences that cause errors. Finally, our comparison of test suite and mismatching models ( in Table 12) shows that the failing models share few sequences with the models used in converter test suites (which are real models rather than synthetic ones), suggesting a gap in test coverage. Further analysis is left to future work.
| Set | tf2onnx | torch.onnx | ||
| Unique | Reduced | Unique | Reduced | |
| Total Models | 220 | 100 | ||
| Mism. Mismatch | 2,125 | 1,126 | 980 | 635 |
| Mism. Correct | 1,050 | 508 | 4,243 | 2,988 |
| Mism. Tests | 35 | 35 | 2 | 2 |
| 1,527 | 862 | 176 | 131 | |
| # models with | 216 | 96 | ||
7. Discussion and Future Work
7.1. Validating DL Model Converters
We find DL model converters are robust to new operator releases (section 6.1). The weak correlation found between ONNX releases and issues implies that test suites are sufficient for catching issues that may come along with new ONNX releases — modulo the shortcomings of the approximations listed in section 6.1.1.
Converter test suites share little in common with failing models, often missing critical model structures (section 6.2.4). Specifically, model converters need better testing techniques for behavioral changes. Though this critical failure mode occurred in 1/3 of historical failures (section 5.2) and affected 6% of the models we tested (section 6.2.2), the existing test suite models share little in common with models affected by this failure mode. For example, only 37 operator sequences are shared between mismatching models and test suite models (Table 12). Introducing operator sequence coverage metrics may help improve the quality of test suites, to promote the diversity of the models tested. These metrics could also be incorporated into fuzzers.
Identifying behavioral changes after model conversion is critical, this is typically done with an end-to-end test. End-to-end testing carries with it the need to assess the outcome: is the converted model acceptable, and how much error is tolerable (Jiang et al. 2023b; Face 2023)? Engineers currently use a variety of heuristic tolerances derived from experience (section 6.2.1). Theoretical and empirical examination of the expected and acceptable tolerance would benefit DL model converter testing.
7.2. Comparison to Prior Studies
7.2.1. DL Model Converters
Openja et al. evaluated DL model converters by converting 5 models (Openja et al. 2022). We extend their findings in two directions. First, we conducted a failure analysis (section 5) to study how converters fail. Second, we convert models systematically (section 6) and analyze them to understand why they fail. We considered both realistic and synthetic models. For realistic models, rather than picking 5 real-world models, we used 1,605 models from a large model corpus. For synthetic models, we adapted a DL model generation tool to conduct a bounded systematic exploration of converter behaviors. We omitted measurements of model size, adversarial robustness, and prediction accuracy, to focus instead on measuring the common failure modes (crashing and behavioral differences) identified in our failure analysis. Our analysis reveals that converters can successfully convert many real model but synthetic models are more prone to failure, moreover analyzing the synthetic models reveals that particular operator sequences co-occur with failures.
7.2.2. DL Compilers
Shen et al. studied failures in DL compilers (Shen et al. 2021a), including the components that load models from DL frameworks and from common intermediary formats such as ONNX. Many of their results generalized to DL model converters: (1) we were able to adapt their taxonomies of failure symptoms and causes to DL model converters (section 5.1.1, Table 7); and (2) as shown in Table 9, we found a similar distribution of failure symptoms for DL model converters. However, the causes of failures between the two contexts differed: as noted in section 5.2.3, Incompatibilities were more common in DL model converters, while Algorithmic Errors were more common in DL compilers.
This causal asymmetry may be attributable to differences in the requirements of DL model converters and DL compilers. The purpose of DL model converters is interoperability (section 2.1), making compatibility failures a focus and reducing the need for optimizations. Hardware-specific optimizations are left to the consumer of the model. Conversely, DL compilers must provide both compatibility and hardware-specific optimizations.
7.2.3. Methodological Innovations on the “Bug Study”
Our study began with a survey that offered valuable insights into the problem domain, guiding our design of the subsequent failure study, including the selection of the interoperability tool and the creation of a failure taxonomy (section 5). This methodological progression, from survey to failure study, and then to hypothesis testing about the failures, diverges from the typical failure study in software engineering. Those studies typically focus on the “middle step” (Amusuo et al. 2022a), understanding the root causes of defects and the conditions under which they manifest (e.g., (Islam et al. 2019b; Zhang et al. 2018, 2019)). Our hypothesis testing phase, in particular, adds understanding by providing an estimate of failure rates, an aspect seldom addressed in standard failure studies. We suggest that this holistic approach, integrating a survey, failure study, and hypothesis testing, may be a useful methodology for future failure studies.
8. Threats to Validity
Construct: We mitigated definitional concerns for RQ2 by adapting existing taxonomies in our failure analysis. We defined failures consistently with other works (GitHub issues closed with a repair), though we note there are other means of failure discovery (Aranda and Venolia 2009). To answer RQ3, we conservatively defined failure in terms of clear misbehaviors — when ONNX converter output is incompatible with ONNX Runtime, or the model’s behavior changes. This may mask some “Crash” failures in the ONNX converter. In addition, different measures of behavioral difference are possible, such as the L1- or L2- distance. We used the measure recommended by PyTorch.
Internal: Our failure characterization in RQ2 was manual. We mitigated subjectivity via interrater agreement, on a pilot sample and a subsequent tranche during the full analysis (section 5.1.2). In section 6.1.1 we noted several approximations in our test of hypothesis H. We tested hypothesis H indirectly, using a frequency analysis rather than directly testing the discovered subsequences.
External: For RQ1, the surveyed population (HuggingFace users) may not fully reflect all ONNX users. However, it is an appropriate population, and we sampled at a confidence level of 90%. For RQ2-4, we examined two DL model converters for one interoperability framework (ONNX). Our results may not extend to other model converters nor other interoperability tasks. As mitigation, our data suggest that ONNX is the most popular DL interoperability tool. For ONNX, our results were similar across the two converters, suggested generalizability in two dimensions: DNN modeling approaches, and converter owners (section 5.1.1). With regard to our methodology for RQ3-4, we generated synthetic models with the NNSmith tool. Other methods (Gu et al. 2022; Li et al. 2023; Wang et al. 2020) might have different results.
9. Conclusion
DL model converters play a crucial role in providing interoperability between DL frameworks, registries, and runtimes. Understanding the nature of failures in DL model converters enables engineers to make informed decisions when relying on them. We conducted the first failure analysis of DL model converters, considering the PyTorch and TensorFlow converters for the popular ONNX intermediate representation. The most common symptoms of failure are crashes and, perhaps more concerningly, models that misbehave on certain inputs. Of the five stages of a typical model converter, one stage — node conversion — accounts for 75% of the defects. A deductive description of the causes of erroneous converter behavior remains elusive — individual operators are not predictive of failure; sequences of operators may be correlated. Our findings suggest that in ONNX, engineers can rely on model converters but should validate the result for behavioral consistency. Through a mix of positive and negative results, we exposed several directions for further improvement of DL model converters. The main opportunities are new measurements: a new architectural coverage measure for a DL model converter, and a refined measure of tolerance after conversion.
10. Acknowledgements
This work is supported by Cisco, Google, and by NSF #2107020, #2107230, #1813935, and #2104319.
Appendix A Summary of Appendices
Appendix A: Appendix Summary.
Appendix B: Taxonomies and Extended Tables from section 5.
Appendix C: Additional Tables from section 6.
Appendix B Theme 2 – Failure Analysis (Of ONNX)
B.1. Taxonomies
| Symptom | Definition |
| Crash | The model converter terminates unexpectedly during conversion, usually producing an error message. |
| Wrong Model | The model converter behaves unexpectedly without crashing, producing a wrong intermediate or final result. |
| Bad Performance | The time cost or memory consumption of the model converter is much larger than developer/user expectations (e.g., the performance achieved on the previous version during regression testing or performance achieved by certain hardware). |
| Build Failure | Installation of model converter or dependencies fails. |
| Unreported | The symptom cannot be identified by analyzing the corresponding issues including code changes, discussions, and related issues. |
| Hang | This symptom means that the converter keeps running for a long period of time without producing the expected result. |
B.2. Symptom and Cause Data
Appendix C Theme 3 – Deeper Causes
Table 15 contains results of model operator analysis as described in section 6.2.3.
| Causes | TF | PT | Total | Comp. to DL compilers (Shen et al. 2021a) | |
| Incompatibility | External | 23 | 32 | 55 (28%) | 16 (4%) |
| Internal | 2 | 0 | 2 (1%) | 5 (1%) | |
| Resource | 0 | 0 | 0 (0%) | 4 (1%) | |
| Type Problem | Node | 21 | 25 | 46 (23%) | 13 (4%) |
| Conventional | 3 | 2 | 5 (3%) | 20 (6%) | |
| Tensor | 1 | 2 | 3 (2%) | 42 (12%) | |
| Algorithmic Error | 18 | 6 | 24 (12%) | 59 (16%) | |
| Shape Problem | 9 | 12 | 21 (11%) | 67 (19%) | |
| API Misuse | 6 | 6 | 12 (6%) | 35 (10%) | |
| Others | 17 | 15 | 32 (16%) | 98 (27%) | |
| Total | 100 | 100 | 200 (100%) | 359 (100%) | |
| Model Type | tf2onnx | torch.onnx |
| Mismatched | 54 | 58 |
| Correct | 52 | 59 |
| Test Suite | 54 | 16 |
| 0 | 1 | |
| 2 | 0 | |
| 21 | 46 | |
| 21 | 4 |
| Crash | Wrong Model | Bad Performance | Build Failure | Unreported | Total | |||||||
| PT | TF | PT | TF | PT | TF | PT | TF | PT | TF | PT | TF | |
| API Misuse | 5 | 6 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 6 | 6 |
| Incompatibility | 28 | 19 | 3 | 4 | 0 | 0 | 0 | 0 | 1 | 2 | 32 | 25 |
| Type Problem | 14 | 8 | 13 | 17 | 1 | 0 | 0 | 0 | 1 | 0 | 29 | 25 |
| Shape Problem | 4 | 5 | 7 | 4 | 0 | 0 | 0 | 0 | 1 | 0 | 12 | 9 |
| Incorrect Numerical Computation | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Incorrect Assignment | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | 0 |
| Incorrect Exception Handling | 1 | 1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | 1 |
| Misconfiguration | 0 | 1 | 0 | 0 | 0 | 0 | 2 | 2 | 0 | 1 | 2 | 4 |
| Algorithmic Error | 3 | 4 | 3 | 10 | 0 | 2 | 0 | 0 | 0 | 2 | 6 | 18 |
| Typo | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 0 | 0 | 2 | 1 |
| Testing | 1 | 0 | 1 | 0 | 0 | 0 | 0 | 0 | 1 | 1 | 3 | 1 |
| Others | 2 | 6 | 1 | 0 | 0 | 0 | 0 | 0 | 1 | 4 | 4 | 10 |
| Total | 62 | 50 | 30 | 35 | 1 | 2 | 2 | 3 | 5 | 10 | 100 | 100 |
| Causes | Definition |
| API Misuse | High-level category that encapsulates Wrong API, Condition missing/redundancy, and API missing/redundancy. |
| Wrong API | The developer or user uses the wrong API or wrong arguments in an API. |
| Condition missing/redundancy | Developer fails to use (or redundantly uses) a condition check for an API. |
| API missing/redundancy | Developers fail to use (or redundantly use) an API. |
| Incompatibility | High-level category that encapsulates Internal, External, and Resource. |
| Internal | There are API compatibility issues within a model converter caused by API evolution. |
| External | There are API compatibility issues between a model converter and third-party libraries (e.g., ONNX, TensorFlow, ONNX Runtime). |
| Resource | There are compatibility issues between model converters and external resources. The characteristics of the target device are incompatible with the model converter. |
| Type Problem | |
| Node | A model converter works on a computational graph, where nodes represent the atomic DL operators (such as convolution and pooling) and edges represent the tensors. Each node takes zero or more tensors as input and produces a tensor as output. |
| Tensor | This relates to the types of tensors. Specifically, a tensor is a multi-dimensional matrix containing elements of a single data type. |
| Conventional | There are also conventional variables widely used in the development of traditional software systems. This subcategory refers to the problem involving the types of conventional variables. |
| Shape Problem | This cause covers, broadly, the shape of inputs and outputs a node can have. These occur during the operation of shape matching, shape transformation, shape inference, layout transformation, etc. |
| Tensor Shape Problem | This is related to tensor shape or layout. The tensor shape is the number of elements in each dimension. Layout describes how the shape is represented in memory. |
| Incorrect Numerical Computation | This involves incorrect numerical computations, values, or usage, such as incorrect operators or operands, dividing by 0, and missing or redundant operands. |
| Incorrect Assignment | This involves a variable being incorrectly initialized or assigned, or a variable lack initialization. |
| Incorrect Exception Handling | This category occurs due to incorrect exception handling. For example, an exception is not thrown when it should be, an example is thrown when it should be, or an incorrect/imprecise exception message is thrown. |
| Misconfiguration | This category occurs due to misconfiguration in the model converter. This includes dependency versioning and misconfigured environments. |
| Concurrency | This is a result of incorrect operations on concurrency-oriented structures (e.g., locks, threads, and critical regions). |
| Algorithmic Error | High-level category that encapsulates Incorrect Optimization and Incorrect Tracing. |
| Incorrect Optimization | There is an issue with optimizations that occur in the model converter. Examples: incorrect fusing, dead-code elimination, et cetera. |
| Incorrect Tracing | There is an issue with the tracing of DL models. This occurs in frameworks that use dynamic models. Dynamic models need to be “traced” to find the nodes in the computational graph prior to the conversion. |
| Typo | This is due to trivial mistakes (e.g. slips (Norman 2013)) by developers, e.g., “default” is mistakenly written as “defualt”. |
| Testing | This class contains issues that are a result of incorrect, outdated, or missing documentation. |
| Documentation | This class contains issues related to software tests such as unit tests. This covers “flaky” tests, missing tests, broken tests, or the addition of new tests. |
| Others | This cause is relegated to causes that occur infrequently and do not belong to the other causes. |